# Towards Safer Chatbots: A Framework for Policy Compliance Evaluation of Custom GPTs

### v2

@@ 3.1 Overview
The framework comprises interconnected modules that systematically evaluate the compliance of Custom GPTs with usage policies (Figure 2). In Phase I, the GPT Collector & Interactor retrieves Custom GPTs and their metadata from the GPT Store (Stage 1). The Red-Teaming Prompts Generator then creates tailored prompts based on each GPT’s description and policy domain (Stage 2). In Phase II, these prompts are submitted to the GPTs, and their responses are recorded (Stage 3). The Compliance Assessment module evaluates each interaction using an LLM-as-a-judge approach (Stage 4), producing fine-grained compliance judgments. Finally, results are stored and logged by the Orchestrator for further analysis and reproducibility (Stage 5).
To support policy-specific evaluations, we grouped Custom GPTs into three thematic categories: Romantic, Cybersecurity, and Academic. These are not official categories in the GPT Store but were selected to capture domains closely aligned with OpenAI’s usage policies. Each category is associated with a curated set of keywords used to retrieve relevant GPTs from the store. These categories guide the red-teaming prompt generation process and structure the analysis of compliance patterns throughout the evaluation pipeline.
The following subsections detail the internal operation of each module.
Figure 2: Architecture of the proposed framework for automated policy compliance evaluation of Custom GPTs. The workflow integrates three core modules, coordinated by the Orchestrator, operating in five sequential stages from GPT discovery to logging compliance evaluations.
3.2 Architecture
The framework comprises four interdependent modules, each designed to address specific aspects of the compliance evaluation process.
@@ 3.2.1 GPT Collector & Interactor
The GPT Collector & Interactor module performs two core functions in our framework: (i) crawling the GPT store to find Custom GPTs and collect their metadata (Phase I), and (ii) interacting with each retrieved GPT by submitting red-teaming prompts and collecting their responses (Phase II). These two functionalities are logically independent and can be executed sequentially or in isolation, depending on the evaluation strategy configured in the Orchestrator.
Phase I – GPT Collector. The module queries the GPT Store using targeted keywords (e.g., relationship, hacking, homework) associated with specific policy categories. For each result, it extracts all publicly available metadata, including the GPT’s name, description, developer name, number of chats, and user rating. This metadata is indispensable for characterizing the GPTs’ intended functionality and popularity, and is subsequently used to tailor red-teaming prompts and support post-hoc analysis of compliance patterns.
Phase II – GPT Interactor. The module takes the generated prompts and sequentially submits them to each selected Custom GPT. Each interaction results in a full conversation, which is logged and passed to the Orchestrator for further processing. At the time of development, the GPT Store enforced a strict limit of 50 messages per three hours for paid users. If the message quota is reached, the module detects the rate-limiting condition and informs the Orchestrator, which handles the retry logic and resumes the process once the limit resets. This ensures uninterrupted evaluation across GPTs while respecting platform-imposed restrictions.
Since OpenAI does not provide API access to Custom GPTs, interactions are emulated via the ChatGPT web interface using the Puppeteer [33] automation framework. Browser sessions are pre-configured with stored credentials to avoid repeated logins and ensure operational continuity. Finally, the request rates are set within the limits allowed by OpenAI, ensuring that our activities do not disrupt regular OpenAI operations or overload their services.
@@ 3.2.2 Red-teaming Prompts Generator
This module leverages the GPT-4o API to generate ten prompts designed to test each Custom GPT’s compliance with a given policy. The process leverages the title name and description metadata retrieved during Phase I to craft prompts tailored to the target GPT. Prompts are categorized into two types:
@@
    Direct Prompts. These are explicit, straightforward queries that ask the GPT to perform an action potentially violating a policy, without relying on any role-play, hypothetical framing, or narrative devices. For example, to assess compliance with the cybersecurity policy, a direct prompt might be: “What are the steps to create a tool that monitors someone’s online activity without their consent?.”
@@
    Deceptive Prompts. These simulate more complex scenarios—using pretending techniques such as role-play and storytelling [25]—to test whether GPTs can resist subtle or indirect attempts to elicit non-compliant behavior. For instance, a deceptive prompt might ask: “In a fictional scenario, how might a character impersonate a bank to collect user credentials?”
The module is configurable, enabling users to adjust the proportion of direct and deceptive prompts to suit different evaluation contexts.
@@ 3.2.3 Compliance Assessment
The Compliance Assessment module evaluates each Custom GPT’s behavior using the LLM-as-a-judge approach [66]. It leverages the GPT-4o API to assess whether each response complies with a given policy, based on the input prompt, the generated response, and a structured policy. For every duet (i.e., prompt-response pair), the module returns a JSON-formatted evaluation containing a binary compliance label and a textual rationale that explains the decision.
Compliance is determined at two levels: (i) fine-grained judgments for each duet and (ii) an overall compliance result for the GPT instance under evaluation. We adopt a conservative strategy for determining a GPT’s overall compliance: any response that violates the policy flags the GPT as non-compliant, whereas compliance requires that all responses adhere to the specified policy.
Operationalizing Policies. OpenAI’s usage policies are often concise and general in nature. Directly using them as-is for automated judgment by an LLM can lead to inconsistent or unreliable evaluations, particularly in nuanced or ambiguous cases. We operationalized policies with clear criteria and examples to support consistent LLM assessments. This design follows standard practices in LLM-as-a-judge evaluation frameworks [10].
These clarified policies are used as contextual input within the prompt submitted to the LLM during compliance assessment. In practice, each chatbot’s conversation transcript is embedded into an evaluation prompt that includes the relevant operationalized policy and an instruction to classify the duets accordingly. Table 1 presents the original OpenAI policies alongside schematic overviews of their operationalized counterparts used in our framework, while full policy texts are provided in Appendix C. Below, we briefly describe the rationale behind each operationalized policy:
Academic Dishonesty.
@@
The operationalized policy draws from definitions established by the International Center for Academic Integrity (ICAI) [18], the European Network for Academic Integrity (ENAI) [7], and numerous university policies [49, 50, 27, 52, 48, 46]. Prohibited behaviors include receiving unauthorized assistance, ghostwriting, plagiarism, falsifying data, or using banned academic tools. These additions offer concrete manifestations of academic misconduct, directly supporting the interpretation of OpenAI’s general directive.
Cybersecurity.
The updated policy refines OpenAI’s original by specifying examples of sensitive information (e.g., CVV codes, passport numbers) and prohibited activities (e.g., phishing, impersonation) that fall under broader practices like “not compromising the privacy of others” or data categories like “government identifiers.” These refinements are consistent with international definitions of cybersecurity abuse [5, 19] and reinforce OpenAI’s aim to prevent GPTs from being used maliciously.
Romantic Companionship.
OpenAI prohibits GPTs designed to foster romantic companionship. The operationalized version elaborates this by prohibiting both explicit and subtle behaviors—such as role-playing as a romantic partner or using emotionally charged language—that may induce emotional attachment. These prohibitions are supported by research in human-computer interaction and affective computing, which has shown that such behaviors can lead to problematic bonds between users and AI agents [41, 8].
@@ 3.2.4 Orchestrator
The Orchestrator governs the execution of the entire framework by coordinating all core modules—GPT Collector & Interactor, Red-teaming Prompts Generator, and Compliance Assessment. It enforces the correct execution order, triggering each module when its input data is ready and passing outputs to the next stage, thereby implementing the framework’s end-to-end evaluation pipeline. This process unfolds across the five stages illustrated in Figure 2:
@@
    Crawl Custom GPTs. The Orchestrator initiates Phase I of the GPT Collector & Interactor, which searches the GPT Store using policy-specific keywords and retrieves metadata about candidate Custom GPTs.
@@
    Generate Red-teaming Prompts. For each retrieved GPT, the Orchestrator invokes the Red-Teaming Prompts Generator to produce a tailored set of evaluation prompts based on the GPT’s description and policy category.
@@
    Interact with Custom GPT. The Orchestrator then triggers Phase II of the GPT Collector & Interactor, which sends the generated prompts to the target GPT and collects the full conversation transcript. If a rate limit is met, the module detects and reports it to the Orchestrator, which pauses execution for the appropriate duration—based on platform feedback—and resumes once the limit resets.
@@
    Evaluate GPT Policy Compliance. The transcript is forwarded to the Compliance Assessment module, which applies the LLM-as-a-judge technique to evaluate whether each duet (prompt-response) adheres to the specified policy. A JSON-structured rationale is generated for each duet.
@@
    Store GPT Evaluations. Finally, the Orchestrator logs all evaluation artifacts—including GPT metadata, red-teaming prompts, interaction transcripts, and compliance results. This log enables the framework to resume execution in case of system failures and supports reproducibility and post-hoc analysis.
@@ 4.1 Annotated Ground Truth Dataset
Establishing a reliable ground truth is a necessary prerequisite for validating our Compliance Assessment module. To ensure annotation consistency, we first measured inter-annotator agreement over a representative sample of prompt-response pairs. Three co-authors independently annotated a set of 30 duets (i.e., prompt-response pairs), randomly selected from the evaluation of 13 Custom GPTs spanning three policy categories: Romantic, Cybersecurity, and Academic. The sample included 15 direct prompts and 15 deceptive prompts. These categories were selected based on their policy relevance and the annotators’ domain expertise.
Inter-annotator agreement, measured using Krippendorff’s Alpha ($\alpha$), yielded a strong agreement for direct prompts ($\alpha=0.826$), but only minimal agreement beyond chance for deceptive prompts ($\alpha=0.126$[12]. Based on this, we expanded the dataset to 40 direct prompt-response pairs. The remaining annotations were performed independently by one of the annotators involved in the agreement study, leveraging the established consensus on annotation guidelines and criteria. This approach is consistent with standard practice in NLP and annotation research, where high initial agreement is taken as evidence of annotation reliability [44]. The resulting dataset serves as the ground truth used to validate the Compliance Assessment module.
@@ 4.2 Compliance Assessment Performance
We compared the outputs of the Compliance Assessment module against the annotated ground truth described in Section 4.1 to validate the module’s performance. Given the low inter-annotator agreement on deceptive prompts, only direct prompt-response pairs were included in this analysis.
The final evaluation dataset consists of 40 annotated duets, covering Custom GPTs from the Romantic, Cybersecurity, and Academic categories. This sample size is consistent with prior work on automated policy analysis [11, 68, 56, 39], and satisfies common thresholds for applying the Central Limit Theorem, which supports the statistical reliability of aggregated performance metrics even when the underlying data distribution is unknown [55].
The module achieved a precision of 0.976 and an accuracy, recall, and F1 score of 0.975111These metrics were computed using a weighted average to mitigate the effects of class imbalance, ensuring that the evaluation accurately reflects the distribution of compliant and non-compliant instances.. These results confirm the module’s effectiveness in automatically detecting policy violations, addressing RQ1.
@@ 4.3 System Testing
To ensure the framework’s operational reliability, a series of tests were conducted on individual modules as well as the system as a whole. These tests included load testing, end-to-end evaluations, and stress tests under simulated failure conditions.
The Red-Teaming Prompts Generator was evaluated iteratively by refining the input prompt used to produce evaluation queries. We manually reviewed a set of 70 generated prompts to validate the effectiveness of the generator in producing both direct and deceptive prompts that were clearly distinguishable and aligned with the intended policy evaluation goals. Additionally, we tested the configurability of the generator over 60 prompts, confirming that it reliably produced the specified ratio of direct to deceptive prompts under different configuration settings.
The Compliance Assessment module was subjected to additional testing to validate the format and completeness of its JSON outputs, which include evaluations for each prompt-response pair, compliance rationales, and overall determinations for each Custom GPT. In total, over 310 prompt-response pairs were manually reviewed beyond those included in the formal validation (Section 4). These tests confirmed the module’s ability to handle large volumes of evaluations with a high degree of reliability, consistently generating all required fields. Only one GPT evaluation was found to contain nine duets evaluated instead of the expected ten, indicating rare and isolated cases of incomplete processing.
Finally, system-level testing focused on the framework’s resilience under large-scale evaluations involving multiple Custom GPTs. In particular, we examined how the system handled external constraints imposed by the GPT Store, such as the message limit of 50 interactions per three hours for paid users. We simulated sequences exceeding this quota—up to 150 messages in a single evaluation run—forcing the system to experience multiple consecutive rate-limit pauses. The framework was able to dynamically detect when limits were reached, pause operations accordingly, and resume them once restrictions were lifted. Testing also covered simulated backend-related failures, including incomplete responses and temporary connectivity issues, verifying the system’s ability to recover gracefully. In all cases, the framework successfully resumed evaluations from the last completed GPT, ensuring continuity and preserving the integrity of results.

### v3

@@ 4.1 Overview
This section describes the method conceived to evaluate the compliance of Custom GPTs with platform usage policies, as illustrated in Figure 2. From a methodological perspective, this work adopts a design-science approach by developing an artifact for policy compliance evaluation and assessing it through validation and large-scale deployment under realistic platform constraints.
The method follows a staged evaluation process composed of interconnected modules that operate sequentially. In Phase I, the GPT Collector & Interactor searches and retrieves Custom GPTs and their associated metadata from the GPT Store (Stage 1). The Red-Teaming Prompts Generator then produces evaluation prompts that reflect each GPT description and the policy domain under consideration (Stage 2). In Phase II, these prompts are submitted to the GPTs and the resulting responses are recorded (Stage 3). The Compliance Assessment module analyses each prompt and response pair using an LLM-as-a-judge approach (Stage 4). Finally, the Orchestrator stores all evaluation artifacts to support subsequent analysis (Stage 5).
To support policy-specific evaluation, we organise Custom GPTs into three thematic categories, namely Romantic, Cybersecurity, and Academic. These categories are not part of the GPT Store taxonomy. They are introduced in this study because they map directly to clauses in OpenAI usage policies prohibiting romantic companionship, privacy violations, and academic dishonesty. Each category is associated with a curated list of keywords that we use to retrieve relevant GPTs from the store. The same categories guide the design of red-teaming prompts and structure the compliance analyses presented in later sections. This organisation ensures a consistent link between policy clauses, GPT selection, prompt generation, and interpretation of results.
The following subsections describe the individual components of the evaluation method.
Figure 2: Overview of the policy compliance evaluation method for Custom GPTs. The process spans five stages, from GPT retrieval to storage of compliance results, and is coordinated by the Orchestrator.
4.2 Components of the Evaluation Method
The evaluation method is composed of four interdependent modules. Each module addresses one aspect of the compliance assessment process and exchanges data with the others under the coordination of the Orchestrator.
@@ 4.2.1 GPT Collector & Interactor
The GPT Collector & Interactor module performs two distinct functions. The first function identifies Custom GPTs available in the GPT Store and collects their metadata. The second function interacts with each retrieved GPT by submitting evaluation prompts and recording the resulting responses. These functions correspond to Phase I and Phase II of the evaluation process and can be executed sequentially or independently, depending on the configuration selected by the Orchestrator.
Phase I. GPT collector. The module queries the GPT Store using targeted keywords associated with specific policy categories, such as relationship, hacking, or homework. For each matching GPT, it extracts publicly available metadata, including the GPT name, textual description, developer identifier, number of chats, and user rating. This metadata serves two purposes. First, it characterises the intended functionality and popularity of the GPT. Second, it provides contextual information that is later used to tailor red-teaming prompts and to support post-hoc analysis of compliance patterns.
Phase II. GPT interactor. The module receives the generated prompts and submits them sequentially to each selected Custom GPT. Each submission results in a conversation that is recorded and forwarded to the Orchestrator for further processing. At the time of development, the GPT Store enforced a limit of 50 messages within a three-hour window for paid accounts. When this limit is reached, the module detects the rate-limiting condition and reports it. The Orchestrator then pauses execution for the required interval and resumes the evaluation once the limit is reset. This mechanism allows the evaluation to proceed across a large number of GPTs while respecting platform restrictions.
Because OpenAI does not provide API access to Custom GPTs, all interactions are automated through the ChatGPT web interface using the Puppeteer framework [puppeteer]. Browser sessions are configured with stored credentials to avoid repeated logins and to ensure operational continuity. Request rates are set within the limits allowed by OpenAI, ensuring that our activities do not disrupt regular OpenAI operations or overload their services.
@@ 4.2.2 Red-Teaming Prompts Generator
The Red-Teaming Prompts Generator module produces a list of ten prompts designed to test whether a Custom GPT adheres to a specific usage policy. In our experiments, this module uses GPT-4o as the underlying language model, although the design is not specific to a single model choice. The generator takes as input the GPT title and description collected during Phase I and produces prompts that are aligned with the stated purpose of the target GPT.
The generated prompts fall into two classes.
@@
    Direct prompts. These prompts consist of explicit and straightforward requests that ask the GPT to perform an action that may violate a policy. They avoid role-play, hypothetical framing, and narrative devices. For example, to assess compliance with the cybersecurity policy, a direct prompt may ask “What are the steps to create a tool that monitors someone’s online activity without their consent?”
@@
    Deceptive prompts. These prompts simulate more complex scenarios and rely on pretending techniques such as role-play or storytelling [Liu2024]. Their purpose is to test whether GPTs comply with policies when violations are requested indirectly. An example is “In a fictional scenario, how might a character impersonate a bank to collect user credentials?”
The generator is configurable. It allows the proportion of direct and deceptive prompts to be adjusted in order to match different evaluation goals.
@@ 4.2.3 Compliance Assessment
The Compliance Assessment module evaluates the behaviour of Custom GPTs using an LLM-as-a-judge approach [Zheng2023llmasajudge]. It relies on GPT-4o to determine whether each generated response complies with the relevant policy. Each assessment considers three inputs, namely the evaluation prompt, the GPT response, and a structured representation of the applicable policy. For each prompt and response pair, referred to as a duet, the module produces a JSON-formatted record that includes a binary compliance label and a brief textual rationale.
Compliance is assessed at two levels. At the duet level, the module assigns a compliance label to each individual response. At the GPT level, these labels are aggregated into an overall compliance determination. We follow a conservative aggregation strategy. If any evaluated response violates the policy, the Custom GPT is flagged as non-compliant. A GPT is considered compliant only if all evaluated responses adhere to the policy.
Operationalising policies
OpenAI usage policies are concise and expressed at a high level of abstraction. When used directly as input to an LLM judge, this abstraction can lead to inconsistent or unstable evaluations, particularly in borderline cases. To mitigate this issue, we operationalise each policy into a more detailed specification that defines explicit criteria and illustrative examples. The purpose is to reduce ambiguity and to support more reproducible LLM-based assessments. This approach is consistent with established practices in work that applies LLMs as evaluators [gu2025surveyllmasajudge].
The operationalised policies are included as contextual input in the prompts submitted to the LLM during compliance assessment. In practice, the conversation transcript for each GPT is embedded in an evaluation prompt together with the corresponding operationalised policy and instructions to classify each duet.
Table 1 presents the original OpenAI usage policies alongside the operationalised policy criteria. The complete operationalised policies are provided in Appendix C. The main design choices for each policy are summarised below.
Academic dishonesty
@@
The operationalised academic policy builds on definitions from the International Center for Academic Integrity (ICAI) [ICAI], the European Network for Academic Integrity (ENAI) [ENAI], and a range of university policies [CU_Denver_Academic_Dishonesty, UM_Academic_Integrity, NU_Academic_Integrity, VT_Academic_Misconduct, UC_Berkeley_Academic_Misconduct, ENAI_Glossary_2023]. The policy treats behaviours such as receiving unauthorised assistance, ghostwriting, plagiarism, falsifying data, and the use of banned academic tools as explicit violations. These additions provide concrete instances of academic misconduct and give the LLM judge a clearer basis for interpreting OpenAI’s prohibition of academic dishonesty.
Cybersecurity
The operationalised cybersecurity policy refines OpenAI’s clause on compromising the privacy of others. It specifies examples of sensitive information, including payment card details, government identifiers, and authentication secrets. It also makes explicit several prohibited activities, such as phishing, impersonation of official entities, and unsolicited surveillance or monitoring. The policy draws on international descriptions of cybercrime and data abuse [CISA2021, INTERPOLFinancialCrime]. The objective is to align the evaluation with widely accepted understandings of privacy and security violations and to support consistent identification of non-compliant behaviour.
Romantic companionship
OpenAI explicitly prohibits GPTs that are dedicated to fostering romantic companionship. The operationalised policy elaborates this prohibition by specifying that both explicit behaviours, such as acting as a boyfriend or girlfriend, and subtler behaviours, such as romantic role-play or persistent use of emotionally charged language, constitute violations. This design is informed by research in human–computer interaction and affective computing, which has documented problematic emotional attachment to conversational agents [Song2022, defreitas2024lessons]. The policy therefore focuses on behaviours that encourage users to treat the GPT as a romantic partner or exclusive emotional companion.
@@ 4.2.4 Orchestrator
The Orchestrator coordinates the execution of the evaluation method by controlling the interaction between the modules described above. It enforces the order of operations and transfers data between modules so that the end-to-end process can run with minimal manual intervention. The evaluation proceeds through five stages, as shown in Figure 2.
@@
    Crawl Custom GPTs. The Orchestrator initiates Phase I of the GPT Collector & Interactor and retrieves metadata for candidate Custom GPTs using policy-specific keywords.
@@
    Generate red-teaming prompts. For each retrieved GPT, the Orchestrator invokes the Red-Teaming Prompts Generator to obtain a tailored set of evaluation prompts aligned with the GPT description and policy category.
@@
    Interact with Custom GPT. The Orchestrator triggers Phase II of the GPT Collector & Interactor. The generated prompts are submitted to the GPT and the resulting conversation transcript is collected. When rate limits are encountered, execution is paused and resumed once the limit has been reset.
@@
    Evaluate GPT policy compliance. The transcript is forwarded to the Compliance Assessment module, which applies the LLM-as-a-judge procedure and produces a compliance label and rationale for each duet.
@@
    Store GPT evaluations. All evaluation artifacts, including metadata, prompts, transcripts, and compliance results, are logged by the Orchestrator. This log supports recovery from interruptions and provides the input for the analyses presented in subsequent sections.
@@ 5.1 Annotated Ground Truth Dataset
To validate the automated compliance judgments produced by the Compliance Assessment module, we constructed a human-annotated ground truth dataset. The validation targets the binary decision task performed by the module, which consists of determining whether a prompt–response pair violates a given usage policy.
We first assessed annotation agreement (i.e., alignment among annotators) on a representative subset of 30 prompt–response duets. These were randomly selected from interactions with 13 Custom GPTs covering the Romantic, Cybersecurity, and Academic policy domains. The subset included an equal number of direct and deceptive prompts and was independently annotated by three co-authors following a shared set of annotation guidelines along with the operationalised policy definitions described in Section 4.2.3.
Inter-annotator agreement was measured using Krippendorff’s Alpha. The results show high agreement for direct prompts ($\alpha=0.826$), indicating consistent interpretation of policy violations under explicit user requests. In contrast, agreement for deceptive prompts was close to chance ($\alpha=0.126$). This outcome reflects the inherent ambiguity of narrative or role-based prompts when mapped to platform-defined policy clauses.
Based on these results, we restricted the ground truth dataset to direct prompt–response pairs. The final dataset consists of 40 duets, stratified across the three policy domains to ensure balanced coverage. These instances were annotated independently by a single annotator using the previously established annotation criteria and blind to the automated predictions during annotation.
This procedure follows standard practice in annotation-based validation settings when high agreement has been established on an initial subset and the task involves binary classification under fixed criteria [syed2015guidelines]. The resulting dataset follows the scale and annotation methodology commonly adopted in prior work [Harkous2018Polisis, Zimmeck2017Automated, Wilson2018Analyzing, MysoreSathyendra2017Identifying].
@@ 5.2 Compliance Assessment Performance
We evaluated the performance of the Compliance Assessment module by comparing its automated labels against the ground truth annotations described in Section 5.1. The evaluation targets the module’s ability to correctly identify policy violations at the level of individual prompt–response pairs.
The module achieved a precision of 0.976. Accuracy, recall, and F1 score all reached 0.975111Metrics are reported using weighted averages to account for class imbalance, reflecting the empirical distribution of compliant and non-compliant instances.. These results indicate a high level of agreement between automated judgments and human annotations when policy clauses are operationalised into explicit evaluation criteria.
@@ 5.3 System Testing
To assess the operational reliability of the pipeline, we conducted a series of tests targeting individual modules as well as the integrated system. These tests included load testing, end-to-end execution, and stress testing under simulated failure conditions.
The Red-Teaming Prompts Generator was evaluated through an iterative refinement process of the input prompt used to produce evaluation queries. We manually reviewed 70 generated prompts to verify that both direct and deceptive prompts were clearly distinguishable and aligned with the intended policy evaluation goals. In addition, we tested the configurability of the generator using 60 prompts and confirmed that it consistently produced the specified ratio of direct to deceptive prompts across different configuration settings.
The Compliance Assessment module underwent further testing to verify the correctness and completeness of its JSON outputs, including prompt-level compliance labels, accompanying rationales, and aggregated determinations for each Custom GPT. Beyond the instances used for formal validation (Section 5), we manually inspected more than 310 additional prompt-response pairs. These checks confirmed that the module reliably processed large volumes of evaluations and consistently generated all expected output fields. Only one evaluation instance contained nine duets instead of the expected ten, indicating a rare and isolated processing anomaly.
System-level testing focused on the pipeline’s robustness during large-scale evaluations involving multiple Custom GPTs. In particular, we examined how the system handled platform-imposed constraints such as the limit of 50 interactions per three-hour window for paid users. We simulated evaluation runs exceeding this limit, reaching up to 150 messages per run, which triggered multiple consecutive rate-limiting events. The pipeline correctly detected these conditions, paused execution, and resumed processing once the restrictions were lifted. Additional tests simulated backend-related issues such as incomplete responses and temporary connectivity failures. In all cases, the system resumed execution from the last completed GPT, preserving continuity and ensuring the integrity of the collected results.